# Ontology-driven Prompt Tuning for LLM-based Task and Motion Planning

### v1

@@ IV-B Semantic Tagging
The semantic tagging process and the identification of tasks and objects from the user input are key components of the proposed system. We use the SpaCy library (https://spacy.io/), which provides advanced capabilities for part-of-speech (POS) tagging and dependency parsing. It takes user input in the form of natural language and extracts the key tasks (verbs) and objects (nouns) by analyzing the grammatical structure of the sentence.
@@
Formally, the input sentence can be represented as a sequence of tokens, i.e., $\{t_{1},t_{2},...,t_{n}\}$. Each token $t_{i}$ is assigned a part-of-speech tag $p_{i}$, where $p_{i}\in\{\text{\small VERB},\text{\small NOUN},\text{\small ADJ},...\}$. The system identifies a task $\alpha$ such that: $\alpha=t_{i}\quad\text{if}\quad p_{i}=\text{\small VERB}\quad\text{and}\quad t_{i}\in\texttt{tasks}.$
@@ IV-C Contextual Inference Module
http://www.example.org/kitchen_ontology#
@@
http://www.example.org/kitchen_ontology#
@@
IV-D Perception and Environmental State State Descriptor
Figure 4: Example scenarios validating the proposed approach: (A) Task: Put bowl, banana, and apple on the plate; (B) Task: Clean table, move sugar box, tomato can, and cracker box to the left table, move plate and cup to the right table.
@@ IV-D Perception and Environmental State State Descriptor
IV-E LLM Task and Motion Planner
The LLM-Task planner takes a generated prompt as input and formulates a symbolic plan to be executed by the robotic system. The symbolic plan generation closely follows the methodology used in the baseline approach [9]. For instance, in a user input prompt such as “put the apple and plate on the tray,” the actual final output prompt feed to the LLM-Task Planner after tuning will be the one shown in Fig. 2 inside the Prompt Generator box, where the yellow text comes from the Contextual Inference Engine and the blue text form the Environment State Descriptor. This prompt will be interpreted by the LLM that generates the following symbolic plan:
@@ IV-E LLM Task and Motion Planner
TABLE I: Comparison of LLM-TAMP and Ontology-Driven LLM-TAMP. TPSR: Task planning success rate, EXESR: Execution success rate, CALLs: LLM Calls
Prompt LLM-TAMP Ontology-Driven LLM-TAMP TPSR % EXESR % # CALLs TPSR # EXESR # # CALLs Put bowl, banana and apple in plate (Task 1.1) 100 100 2.8 100 100 3.2 Put banana, apple and bowl in plate (Task 1.2) 60 60 8.6 100 100 2.6 Clean table, move sugar_box, tomato_can, and cracker_box to the left table, move plate and cup to the right table (Task 2.1) 100 90 7.7 100 90 6.3 Clean table, move plate and cup to the right table, move sugar_box, tomato_can, and cracker_box to the left table (Task 2.2) 40 30 8.4 100 90 6.8 Clean table, move crockery items to the right table, move boxed food items to the left table (Task 2.3) 0 0 10 50 40 8.3
@@ V Results and Discussion
The proposed ontology-driven LLM-TAMP framework is validated through simulations as shown in Fig. 4 and real-world experiments depicted in Fig 5, ensuring its applicability and robustness in different task scenarios. We utilized a PyBullet-based simulation environment to test the scenarios. In the first scenario (Fig. 4-A), the objective was to organize items by placing an apple, a banana, and a bowl onto a plate. In the second scenario (Fig. 4-B), the task involved cleaning a table by moving objects. Crockery items such as a cup and a plate are to be moved to the left_table, while boxed food items, including a cracker_box, sugar_box, and tomato_can, to the right_table. These scenarios were designed to evaluate the system’s ability to handle complex object arrangements where the correct temporal order of the subtasks is crucial for performing the task semantically correctly.
@@
Figure 6: Example of a wrong symbolic plan computation.
@@
We conducted a comprehensive evaluation to compare the proposed framework with the baseline approach. The results presented in Table I highlight the importance of the integration of ontology-based reasoning into LLM-TAMP systems. The parameters we compare are task planning success rate (TPSR), execution success rate (EXESR), and the number of LLM calls (# CALLs). The ontology-driven approach consistently outperforms the baseline LLM-TAMP, particularly in complex scenarios that require refined reasoning and object categorization.
For simple tasks where the task is clearly represented in a proper order of objects, such as Put bowl, banana, and apple in plate, the ontology-driven LLM-TAMP and the traditional LLM-TAMP achieve the same TPSR and EXESR of 100%, requiring only a few LLM calls on average. In the same task, if the order of the objects is inverted, the performance of the traditional LLM-TAMP is reduced with the TPSR and EXESR of 60% and a high number of LLM calls, 8.6 on average. Whereas ontology-driven LLM-TAMP maintains the same TPSR, EXESR, and lower call to LLM, 2.6 on average. The reason for better results is that the ontology-driven approach uses knowledge-based reasoning to explicitly specify the correct sequence of object manipulation for LLM. In contrast, the traditional LLM-TAMP approach leaves it to LLM to decide which considers the similar order to manipulate objects as described in the input prompt, which is semantically incorrect, as illustrated in Fig. 6 (put apple and banana first, then place bowl on them). These results confirm the framework’s capability to enhance task and motion planning using knowledge-based reasoning for improved prompt generation and execution accuracy.
In more complex scenarios, but with proper instruction and proper order of objects, such as Clean table, move sugar_box, tomato_can, and cracker_box to the left table, move plate and cup to the right table, the LLM-TAMP and the ontology-driven approach maintain a TPSR of 100% and EXESR of 90% while requiring similar number of calls to LLM. However, in the task if the prompt is given as Clean table, move plate and cup to the right table, move sugar_box, tomato_can, and cracker_box to the left table, the performance of LLM-TAMP is significantly reduced. At the same time, the ontology-driven LLM-TAMP maintains a performance similar to that of the first case. If we provide the prompt in a more generic way, such as Clean table, move crockery items to the correct table, move the boxed food items to the left table the LLM-TAMP fails (0% TPSR and EXESR). The ontology-driven system achieves modest success (50% TPSR and 50% EXESR). The reason for the ontology- driven LLM-TAMP to success is the environment state description using ontological knowledge in which the type of each object is specified by querying on the ontological knowledge (such as cracker_box is a boxed food and plate is a crockery item). This reflects that even in the case of under-specified or ambiguous prompts, where additional context is required, the ontology-driven LLM-TAMP is still able to achieve some success.
In general, ontology-driven enhancements demonstrate improved planning success rates and efficiency across all but the most ambiguous scenarios. Using structured domain knowledge, the system reduces reliance on LLM calls and achieves more consistent and interpretable results, leading to more reliable task planning and execution in complex environments.
Figure 7: Histogram comparing the time required for prompt construction and task planning success rates between LLM-TAMP and ontology-driven LLM-TAMP.
Querying the ontology for prompt tuning required additional time to construct the final prompt for the LLM. To evaluate computational efficiency, we compared the average time taken by LLM-TAMP and ontology-driven LLM-TAMP for prompt construction (Fig. 7). In addition, we analyzed their corresponding task planning success rates. As illustrated in Fig. 7, although ontology-driven LLM-TAMP requires more time for prompt generation compared to standard LLM-TAMP, it consistently outperforms the latter in generating semantically correct symbolic plans. This shows that the improved quality of planning justifies the additional computational overhead in most cases.
VI Conclusion

### v2

@@ IV-B Semantic Tagging
The proposed system relies significantly on semantic tagging and recognizing tasks and objects from user input. We use the SpaCy library (https://spacy.io/), which provides advanced capabilities for part-of-speech (POS) tagging and dependency parsing. It takes user input in the form of natural language and extracts the key tasks (verbs) and objects (nouns) by analyzing the grammatical structure of the sentence.
@@
Formally, the input sentence can be represented as a sequence of tokens, i.e., $\{t_{1},t_{2},...,t_{n}\}$. Each token $t_{i}$ is assigned a part-of-speech tag $p_{i}$.
Where $p_{i}\in\{\text{\small VERB},\text{\small NOUN},\text{\small ADJ},...\}$. The system identifies a task $\alpha$ such that: $\alpha=t_{i}\quad\text{if}\quad p_{i}=\text{\small VERB}\quad\text{and}\quad t_{i}\in\texttt{tasks}.$
@@ IV-C Contextual Inference Module
<http://.../kitchen_ontology#>
@@
http://.../kitchen_ontology#>
@@
IV-D Perception and Env. State Descriptor
@@
Figure 4: Illustrate how the prompt generator integrates the information of user input (green), contextual inference module (yellow), Environment State Description (blue), and Prompt Template (black), to construct the final system prompt. The Generated Prompt incorporates structured environment data, action constraints, and reasoning to guide robotic decision-making effectively.
IV-E Prompt Generator
This module generates the prompt to be sent to the LLM-TAMP module by combining the prompt template with the information coming from the Contextual Inference Engine and Environmental State Descriptor modules. For instance, in a user input prompt such as “put the apple and bowl on the tray”, the actual final output prompt feed to the LLM-Task Planner after tuning will be the one shown in Fig. 4 inside the prompt Generated box, where the prompt template is completed with the yellow text that comes from the Contextual Inference Engine, the blue text that comes from the Environment State Descriptor, and the green text that comes from the user input.
IV-F LLM Task and Motion Planner
The LLM-Task planner takes a generated prompt as input and formulates a symbolic plan to be executed by the robotic system. The symbolic plan generation closely follows the methodology used in the baseline approach [9]. For instance, using the prompt commented in the previous section, the LLM generates the following symbolic plan:
@@ IV-F LLM Task and Motion Planner
Figure 5: Example scenarios validating the proposed approach: the first row shows the initial states and the second row shows the goal state. each scenario is used to perform multiple tasks, some of them are given below: (A) Task: Put bowl, banana, and apple on the plate; (B) Task: Clean table, move sugar box, tomato can, and cracker box to the left table, move the plate and cup to the right table; (C) Task: Serve breakfast by placing plate, bread, and cup on the table; (D) Task: Stack plate1, plate2, and cup on plate3.
@@ V Results and Discussion
This section provides a comprehensive evaluation of the performance of the Onto-LLM-TAMP framework and its critical components, including semantic tagging accuracy, effectiveness of knowledge-based reasoning, and comparative benchmarking of various LLM models.
V-A Experimental Setup
Figure 6: Sequence of snapshots of the following tasks: (A) Task: Put apple, banana, and bowl in plate; (B) Task: Clean table, move plate and cup to the right_table, move sugar_box, tomato_can, and cracker_box to the left_table.
The proposed Onto-LLM-TAMP framework was comprehensively validated through both simulation and real-world experiments, confirming its practical applicability and robustness across diverse task scenarios. Fig. 5 illustrates the simulation scenarios used for validation. These scenarios were implemented using the PyBullet physics engine. Four different task scenarios were designed to thoroughly evaluate the system’s performance.
In the first scenario (Fig. 5-A), the task objective involved organizing items by placing an apple, a banana, and a bowl onto a plate. The second scenario (Fig. 5-B) focused on a table-cleaning task that required the relocation of various objects; specifically, crockery items (cup and plate) had to be moved to the left_table, whereas boxed food items (cracker_box, sugar_box, and tomato_can) were to be placed on the right_table. The sequence of execution is depicted in Fig. 6.
@@ V-A Experimental Setup
Figure 7: Example of a wrong symbolic plan computation. for the Task: Put apple, banana, and bowl in plate.
The third scenario (Fig. 5-C) simulated arranging a breakfast table, involves the precise placement of a cup and plate on the table, followed by putting a bread on the plate. The final scenario (Fig. 5-D) is about stacking objects, where (plate1, plate2, and cup) had to be systematically stacked onto plate3. To further assess the system’s robustness, variations in task prompts were introduced by changing the order of objects mentioned by the user. These scenarios were carefully designed to evaluate the system’s capability in managing complex arrangements of objects, where maintaining the correct temporal sequence of subtasks is critical for achieving semantically correct and successful task execution.
TABLE I: Comparison of LLM-TAMP and Onto-LLM-TAMP. TPSR: Task planning success rate, EXESR: Execution success rate, CALLs: LLM Calls
Prompt LLM-TAMP Onto-LLM-TAMP TPSR % EXESR % # CALLs TPSR # EXESR # # CALLs Task 1: Put bowl, banana and apple in plate 99.5 97.3 2.8 99.7 97.1 3.2 Task 2: Put banana, apple and bowl in plate 30.7 29.4 8.6 99.4 96.9 2.6 Task 3: Clean table, move sugar_box, tomato_can, and cracker_box to the left table, move plate and cup to the right table 96.4 89.3 7.7 98.8 93.2 6.3 Task 4: Clean table, move plate and cup to the right table, move sugar_box, tomato_can, and cracker_box to the left table 42.8 36.6 8.4 98.3 91.4 6.8 Task 5: Clean table, move crockery items to the right table, move boxed food items to the left table 0 0 10 48.1 41.6 8.3 Task 6: Serve breakfast by placing plate, bread and cup on the table. 98.4 90.3 4.6 98.7 90.1 4.9 Task 7: Serve breakfast by placing bread, plate and cup on the table. 20.8 19.6 9.7 98.6 92.4 4.3 Task 8: Stack plate1, plate2 and cup on plate3 98.9 94.8 6.7 98.7 92.1 5.3 Task 9: Stack cup, plate1, and plate2 on plate3 17.2 12.4 9.8 98.1 93.2 5.8 Task 10: Stack cup and plates on the table 0 0 10 68.1 61.8 6.3
V-B Onto-LLM-TAMP Evaluation
@@
The above scenarios are used to compare the proposed framework with the baseline approach. The results presented in Table I highlight the importance of integrating ontology-based reasoning into LLM-TAMP systems. The parameters we compare are;
  • Task planning success rate (TPSR): Represents the percentage of tasks for which the generated plans by the LLM correctly capture the intended task semantics and objectives, resulting in an executable action sequence.
  • Execution success rate (EXESR): Indicates the percentage of tasks successfully completed when executing the generated plans on the robot, highlighting the practical feasibility and effectiveness of the plans.
  • Number of LLM calls (# CALLs): Measures how many LLM calls were required to generate a suitable action plan, reflecting the efficiency and complexity involved in the planning process.
The Onto-LLM-TAMP consistently outperforms the baseline LLM-TAMP, particularly in complex scenarios that require refined reasoning and object categorization. For straightforward tasks with explicit and correctly ordered object descriptions, such as Task 1: Put bowl, banana, and apple in plate, both Onto-LLM-TAMP and traditional LLM-TAMP systems achieve nearly perfect performance (approximately 99.5% TPSR and above 97% EXESR), requiring only a minimal number of LLM calls. However, when the order of objects is altered, as seen in Task 2, the traditional LLM-TAMP’s performance significantly drops to about 30.7% TPSR and 29.4% EXESR, along with an increased number of LLM calls (8.6 on average). In contrast, Onto-LLM-TAMP maintains high performance (99.4% TPSR and 96.9% EXESR) with fewer LLM calls (2.6 on average). This robustness improvement for the Onto-LLM-TAMP is due to the explicit incorporation of the correct semantic sequence of actions using knowledge-based reasoning, which reduces ambiguity in the task. The reason for the failure of LLM-TAMP is that in most of the cases, it ended up performing the task as shown in Fig. 7.
In more complex scenarios, such as Task 3, which involves detailed instructions about moving various items (crockery and boxed foods) to specific locations, both approaches initially perform similarly with high success rates (over 96% TPSR and around 90% EXESR). However, when the task description sequence is reversed and made ambiguous (as in Task 4 and Task 5), the baseline approach suffers a significant performance decline, reaching as low as 0% in TPSR and EXESR. The Onto-LLM-TAMP, meanwhile, maintains a consistently high-performance level (TPSR around 98% and EXESR around 91% for Task 4, and modest success of approximately 48% TPSR for the most ambiguous Task 5). This capability results from using structured ontological knowledge to precisely categorize items (e.g., identifying cracker_box as boxed food and plate as crockery), thus providing the LLM with clearer contextual information.
For tasks that involved serving and stacking items (Tasks 6 to 10), the Onto-LLM-TAMP consistently demonstrates superior performance compared to the baseline. In Tasks 6 and 7, which involve serving breakfast, the baseline’s TPSR drops to around 20% when the item order is changed. In contrast, the Onto-LLM-TAMP maintains consistently high success rates (above 98% TPSR and around 92% EXESR). Similarly, in stacking tasks (Tasks 8 and 9), the Onto-LLM-TAMP achieves high TPSR (around 98%) and EXESR (above 92%), significantly outperforming the baseline, especially when the item order is varied. Task 10, involving stacking items ambiguously described as cup and plates on the table, highlights the limits of both approaches; the Onto-LLM-TAMP approach still achieves success (68.1% TPSR and 61.8% EXESR) because the knowledge reasoner described: the objects with the large bounding boxes will come first in stacking. These results indicate the Onto-LLM-TAMP approach’s capability to handle ambiguity and complex semantic scenarios more effectively.
Figure 9: Screenshots of pose estimation during the execution in the real environment. The task was to put banana, apple, and bowl into the plate.
During the real-world demonstration, we execute the task of put a banana, an apple, and a bowl in the plate, showcasing the system’s ability to accurately detect, localize, and manipulate objects in a real environment. Figure 8 presents a series of screenshots capturing key moments of the execution process. To ensure precise object identification and spatial positioning, we utilize YOLO-V8 for object detection and FoundationPose for pose estimation. YOLO-V8 enables real-time recognition of objects in the scene, while FoundationPose estimates their positions and orientations, allowing the robot to plan and execute accurate pick-and-place actions. Figures 10 and 9 illustrate the intermediate steps, including detected objects and their estimated poses, demonstrating the effectiveness of our approach in real-world conditions.
V-C Semantic Tagging and Reasoning Evaluation
TABLE II: Processing time and accuracy of semantic tagging (ST) and reasoning process (RP) while tuning the prompts.
Experiment Time (s) Accuracy (%)
ST RP ST RP
Exp 1 (Task 1-10) 0.0048 0.094 98.43 89.71
Exp 2 (Random Tasks) 0.0071 0.109 81.51 77.94
Semantic tagging and knowledge-based reasoning are two core components of the proposed framework. Semantic tagging analyzes the input prompt to identify the intended action and associated objects. Subsequently, the knowledge-based reasoner utilizes these extracted actions and objects to enrich and clarify the input, resulting in a more precise and effective task description. Table II presents two distinct measures for evaluating the performance of semantic tagging and knowledge-based reasoning.
Figure 10: Screenshots of object identification during the execution in the real environment. The task was to put banana, apple, and bowl into the plate.
  • Time: It quantifies how long it takes for semantic tagging to extract actions and objects, as well as how long the ontology-based reasoner needs to process those extracted elements.
  • Performance accuracy: Indicates how precisely each component performs its function: For semantic tagging, the percentage reflects the accuracy with which actions and objects are identified in the input prompt. For reasoning, it shows how accurately the ontology-based approach infers relationships and enhances the user prompt.
Table II uses these measures to highlight the speed and precision of the two main processing steps. Overall, the results show that semantic tagging is consistently fast yet experiences a drop in accuracy when dealing with less structured input. Meanwhile, reasoning involves slight computational overhead to perform ontology-driven inferences and similarly shows fairly good accuracy when dealing with a more diverse language, reflecting the impact of varying input complexity.
In experiment 1, we applied semantic tagging and reasoning on each task 10 times and computed the average time and accuracy for 10 tasks. Semantic tagging requires an average of 0.0048 seconds to process the input string and achieve an accuracy of 98.43%. This suggests that for these structured tasks, the tagging process is efficient and highly accurate in recognizing relevant actions and objects. Meanwhile, the reasoning step reports an average processing time of 0.094 seconds with an accuracy of 89.71%. This indicates that once actions and objects are identified, the reasoner uses the ontology to derive correct inferences and enhance the input prompt most of the time. However, a small fraction of tasks may involve entity mismatches or incomplete definitions.
In experiment 2, we used random task prompts. Comparable to tasks 1-10, but introduced some typos, misplaced commas, and other parts of speech. The semantic tagging time increases slightly to 0.0071 seconds. It reflects that tasks expressed in less predictable language can require slightly more processing. The corresponding accuracy of 81.51% shows that the NLP component finds the correct actions and objects in most but not all cases due to more variable and ambiguous language. The reasoner’s processing time is 0.109 seconds, indicating that it can handle these more diverse inputs but does so with a bit more computational load. Its accuracy of 77.94% suggests that knowledge-based inference remains generally effective but can occasionally fail to match extracted entities to the ontology when the input is unfamiliar.
TABLE III: ”Comparison of LLM models on Tasks 1 to 10 for LLM-TAMP and Onto-LLM-TAMP, evaluated based on two key metrics: execution time (T) in seconds and correctness (C) score in %.
Task GPT-4 Gemini Cohere LLaMA
LLM-TAMP Onto-LLM-TAMP LLM-TAMP Onto-LLM-TAMP LLM-TAMP Onto-LLM-TAMP LLM-TAMP Onto-LLM-TAMP
T C T C T C T C T C T C T C T C
1 2.55 100 2.62 100 2.31 100 2.15 100 3.01 100 3.03 100 3.92 100 3.98 100
2 3.25 0 3.61 100 2.10 0 2.41 100 3.02 10 2.44 100 3.57 0 3.32 100
3 3.19 100 4.87 100 3.47 100 2.96 100 3.07 30 3.35 30 4.91 90 5.28 90
4 2.86 10 3.66 100 3.53 10 3.48 100 3.06 0 3.49 0 3.75 10 2.89 100
5 3.12 10 3.80 100 3.53 10 3.48 100 3.61 0 2.90 0 4.61 10 4.96 50
6 3.23 100 3.72 100 1.99 100 2.71 100 3.75 90 3.24 90 4.75 90 4.45 90
7 2.60 10 3.70 100 2.68 10 2.74 100 4.23 10 3.31 90 5.09 0 4.95 90
8 3.60 100 3.62 100 2.93 100 2.06 10 3.18 100 3.92 100 3.22 10 4.91 10
9 3.82 10 3.72 100 2.50 10 2.86 10 4.01 10 3.28 10 4.03 10 4.23 10
10 4.29 10 4.65 10 2.86 10 2.74 10 3.35 10 3.12 10 4.04 10 4.80 100
Average 3.25 55.0 3.80 91.0 2.79 55.0 2.76 83.0 3.43 36.0 3.21 63.0 4.19 34.0 4.28 74.0
Figure 11: Histogram showing the average correctness scores of LLM models for LLM-TAMP and Onto-LLM-TAMP cases.
V-D LLM Models Evaluations
The quality and correctness of the computed symbolic plan largely depend on the clarity of the prompt and the capabilities of the LLM model. To evaluate this, we assessed the performance of various LLM models, specifically GPT, LLaMA, Gemini, and Cohere in both standard LLM-TAMP and onto-LLM-TAMP frameworks. The results of this evaluation are summarized in Table III. The experiments measure two key parameters for both approaches:
  • Time: Defined as the duration each LLM requires to generate executable task plans using API call.
  • Correctness:, Assessed by two robotics experts who manually evaluated each generated action sequence. The evaluation criteria combined semantic accuracy (90%) and structural correctness (10%). High correctness scores indicate that the generated plans are not only syntactically executable but also semantically coherent and realistically applicable to robotic scenarios.
Figure 12: Histogram showing the average API call time of LLM models for LLM-TAMP and Onto-LLM-TAMP cases.
For tasks with straightforward, clearly structured descriptions (e.g., Task 1 and Task 8), both methods performed comparably well, achieving high correctness scores (ranging from approximately 92% to 100%). This consistency demonstrates that, when task instructions are explicit and sequentially correct, LLM-based approaches reliably translate user prompts into executable task plans. Time performance in such tasks varied slightly while generally remaining stable across models and methods. However, when tasks contained ambiguity, non-explicit ordering, or object arrangements different from the correct execution sequence (e.g., Task 2: ”Put banana, apple, and bowl in plate,” Task 4, and Task 5), the performance of the LLM models significantly declined. It can be observed by correctness scores that drop sharply (often to 0–10%) when the user task prompt is directly provided as input to the LLM model (LLM-TAMP). Such tasks posed a major challenge for LLM-TAMP due to the inherent ambiguity or lack of structured context in the provided input, leading to an unreliable sequence of actions. In contrast, when the user prompts along with the ontological reasoning passed to the LLM Models (onto-LLM-TAMP), it consistently maintained high correctness (usually around 100%), clearly showcasing its ability to handle ambiguity through ontological knowledge integration effectively. By utilizing structured domain knowledge of onto-LLM-TAMP, the LLM models were able to interpret ambiguous instructions more precisely, ensuring semantically coherent action sequences.
Interestingly, even though onto-LLM-TAMP involves an additional reasoning step that potentially increases the overall processing complexity, it still achieves comparable execution times in most of the cases. This indicates that ontology-driven clarification of task semantics can streamline the ensuing LLM processing by providing clearer input.
The average results presented at the end of Table III clearly illustrate the overall advantages of employing an Onto-LLM-TAMP approach compared to the standard LLM-TAMP setting across all evaluated models. Specifically, the average correctness improved significantly when using the ontological enhancement, reflecting enhanced semantic understanding and reduced ambiguity in generated action sequences. For clarity, we have added a histogram showing the average correctness score of LLM models for the cases of LLM-TAMP and Onto-LLM-TAMP. In correctness, GPT-4 showed an improvement from 55% correctness in standard LLM-TAMP to 91% correctness with ontology integration. Similar trends are observed for Gemini (from 55% to 83%), Cohere (from 45% to 63%), and LLaMA (from 34% to 74%).
In terms of computational efficiency, measured by the average API calls response time, the Onto-LLM-TAMP maintained comparable response times compared to the basic LLM-TAMP method. Fig. 12 shows the histogram of the comparison of average API call time of LLM models in LLM-TAMP and Onto-LLM-TAMP. For GPT-4 and LLaMA, the Onto-LLM-TAMP required a modest increase in average time (3.25s to 3.80s and 4.19s to 4.28s, respectively). Conversely, Gemini and Cohere remained stable (Gemini: 2.79s to 2.76s; Cohere: 3.43s to 3.21s). These variations indicate that while ontology integration slightly affects response times, the substantial improvements in correctness significantly surpass minor time trade-offs.
V-E Motion Planning Time Evaluation
TABLE IV: Average Motion Planning Time (seconds) for Tasks 1–10, Comparing LLM-TAMP and Onto-LLM-TAMP.
Task Motion Planning Time (s)
LLM-TAMP Onto-LLM-TAMP
Task 1 0.13 0.11
Task 2 2.62 0.14
Task 3 0.61 0.60
Task 4 6.48 0.93
Task 5 NA 3.72
Task 6 0.33 0.28
Task 7 1.89 0.30
Task 8 0.27 0.25
Task 9 4.36 0.23
Task 10 NA 1.59
Motion planning is another critical component of the proposed framework. Its performance significantly influences the overall efficiency of task execution. To evaluate this aspect, we compared the motion planning performance of the standard LLM-TAMP and the Onto-LLM-TAMP approaches using the RRTConnect algorithm. The results are summarized in Table IV. Motion planning times were calculated as the average planning time of multiple calls of the motion planning during the execution of the subtask of each given task.
For relatively simple tasks (Tasks 1, 3, 6, and 8), both approaches exhibited comparable planning times, indicating the minimal impact of ontological reasoning when task instructions are clear and straightforward. For instance, Task 1 demonstrated similar results (0.13s vs. 0.11s), while slightly more complex tasks, like Task 3, showed nearly identical performance (0.61s vs. 0.60s). However, even moderately complex tasks with slight ambiguity, such as Task 2 and Task 7, significantly benefited from ontology-driven reasoning, resulting in notably reduced planning times (from 2.62s to 0.14s and 1.89s to 0.30s, respectively). Ontology-driven reasoning showed significant advantages in complex or ambiguous scenarios (Tasks 4, 5, 9, and 10). For example, in the case of Task 4, planning times were reduced from 6.48s to 0.93s. Tasks 5 and 10, completely failed under standard LLM-TAMP, and succeeded efficiently (3.72 and 1.59s, respectively) with the Onto-LLM-TAMP.
The primary reason for the significant reduction in planning time observed with the Onto-LLM-TAMP is that, in scenarios where task instructions or object ordering are ambiguous or incorrect, the conventional LLM-TAMP method often struggles to find collision-free trajectories. This struggle causes repeated failures at the motion planning stage, which then prompts the LLM planner to recompute the symbolic plan continuously, consequently increasing the total motion planning duration, particularly in complex tasks.
VI Conclusions